A general number m varies directly as another general number n. When n=-6 1/4, m=-7 1/3. Find the value of m when n=2. Thanks so much!
Found 2 solutions by stanbon, DrBeeee:
Answer by stanbon(75887)

A general number m varies directly as another general number n.
m = k*n
-------
Solve for "k" using "When n=-6 1/4, m=-7 1/3.
-7 1/3 = k*(-6)
-22/3 = -6k
k = 11/9
-------
Equation:
m = (11/9)n
Find the value of m when n=2.
m = (11/9)2
----
m = 22/9 = 2 4/9
======================
Cheers,
Stan H.
================

Answer by DrBeeee(684)

The equation for a direct variation is
(1) m = c*n, where c is a constant
It is the same as a straight line with a slope of c and a y-intercept (m-intercept here) of zero. That is, it goes through the origin as (1) does. When n=0 we have
(2) m = c*0 or
(3) m = 0
We need only one point on the line to find the constant c. This is
(4) (n.m) = (-25/4,-22/3)
Put these coordinants into (1) and get
(5) -22/3 = c*(-25/4) or
(6) c = (-22/3)*(-4/25) or
(7) c = 88/75, giving the equation
(8) m = 88/75*n
Now when n=2 we get
(9) m = (88/75)*2 or
(10) m = 176/75
Answer: when n=2, m=176/75